# Multi-Source EEG Emotion Recognition via Dynamic Contrastive Domain Adaptation

### v1

@@ VII Conclusion
With the use of frequency information of EEG signals, our MS-DCDA model demonstrated improved emotion recognition performance. The spatial features of the EEG data have not yet been fully explored due to the lack of spatial information of the electrodes in the current dataset. Exploring the spatial information of electrodes without significantly increasing the collection cost and investigating multi-modality learning that combines frequency and spatial EEG features warrants a future study. Future work may involve extending the examination of EEG features in the time domain. The accuracies, generalization, and robustness of our model may also benefit from exploring the attention mechanism for MLPs.

### v2

@@ 6 Discussion
While our MS-DCDA model demonstrated improved emotion recognition performance with the use of frequency information of EEG signals, the accuracies of the emotion recognition are found to be significantly reduced without the use of DE features regardless of the recognition models. This observation reaffirms the advantage of DE features as an effective representation of EEG signals for emotion recognition. Future work may benefit from more advanced and learnable feature representations based on neural networks. End-to-end learning of EEG representation with subsequent emotion recognition models may further improve accuracy. In addition, the spatial features of the EEG data have not yet been fully explored due to the lack of spatial information of the electrodes in the current dataset. Collecting the spatial information of electrodes without significantly increasing the experiment cost may require further investigation. Multi-modality learning that combines frequency and spatial EEG features also warrants a future study. Future work may involve extending the examination of EEG features in the time domain. The accuracies, generalization, and robustness of our model may also benefit from exploring the attention and gating mechanisms.
Figure 5: Comparison of emotion recognition accuracies of different domain adaptation algorithms on SEED and SEED-IV datasets respectively for (a) Cross-subject experiment; and (b) Cross-session experiment. Our MS-DCDA model outperforms in all experiments four single-source methods: DCORAL [58], DAN [41], DANN [42], DDC [57]. The performance of our model also exceeds two multi-source methods: MS-MDA [56], and MS-ADA [50].
7 Conclusion
In summary, a multi-source dynamic contrastive domain adaptation (MS-DCDA) method was introduced for EEG based emotion recognition. The proposed combines coarse- and fine-grained domain alignments and dynamically optimize the domain transferability and discriminability, leading to improved model generalization, classification accuracies, wider inter-class margin, and higher intra-class compactness. The proposed MS-DCDA outperformed both the classical and the state-of-the-art methods on both the SEED and SEED-IV datasets by a large margin. In addition, Our study also suggests greater emotional sensitivity in the frontal and parietal brain lobes, providing insights for mental health care. The broader significance of this research also lies in its potential applications in human-computer and brain-computer interfaces, allowing systems to respond more precisely to the emotional states of users through enhanced emotion recognition capabilities.
Acknowledgments
This work is supported by the National Natural Science Foundation of China Grants 62372371 and 61972315.